OpenStudy (anonymous):

Can anyone help explain how to write an algebraic expression from a word problem with percentages?

OpenStudy (anonymous):

Sure. Post the problem.

OpenStudy (anonymous):

The final grade for this course is calculated by dividing total points earned by total points available, and a grade of 72% is required to earn a C. There are 618 points available in the first four weeks of the course and 255 points in the last two weeks of the course. If a student earns a grade of 63% in the first four weeks of the course, what percentage grade must the student earn in the last two week in order to earn a final grade of C? She will accept an answer that is expressed as a decimal.

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

In the first four weeks of the course the student has earned 0.63*618=389 points. To earn a C in the course a student needs 0.72*(618+255)=629 points at the end of 6 weeks. The student has already earned 389 points and so needs another 629-389=240 points. There are 255 points available and the student needs to earn 240 of them So the student needs 240/255=0.94 or 94% in the last two weeks to get a C in the course.

OpenStudy (anonymous):

i understand that for the most part, but she wants us to use 1 variable in writing the equation, how would you take that into account?

OpenStudy (anonymous):

ok. I did it this way to kind of guide you through it but it can be written more succinctly. Hang on...

OpenStudy (anonymous):

Okay. So we write and expression for total points earned divided by total points available and se that equal to the grade we want, i.e., 0.72:\[\frac{(0.63)(618)+255x}{873}=0.72\] Solve for x.